OpenStudy (anonymous):

Cut points of (x-2)(x+2)(3x-1)≤0: a. 2, -2, 3 b. 2, -2, -1/3 c. 2, -2, 1/3 d. 2, -2, -3

OpenStudy (anonymous):

Well, do you know what valeus of x will make (x-2)(x+2)(3x-1) 0?

OpenStudy (anonymous):

(x-2)(x+2)(3x-1) = 0 is what I mean. What x values make that true?

OpenStudy (anonymous):

Is the answer c.? I just need clarification

OpenStudy (anonymous):

Yes. But you probably didn't need to use Alpha. If you look at (x-2)(x+2)(3x-1) (x-2) = 0 gives me x=2 (x+2) = 0 gives me x=-2 (3x-1) = 0 gives me x=(1/3)

OpenStudy (anonymous):

Oh I see. Thank you for the explanation and the clarification!

OpenStudy (anonymous):

You're quite welcome =)